• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dutta

    Log entry time 06:03:42 on February12,2002

    Entry number 82671

    keyword=shift summary

    Shift Summary
    Owl shift 2/12/2002
    M. Roedelbronn, E. Chudakov, D.Dutta

    Beam back at 00:15
    00:16 start run 2468 kine 5G-
    00:30 Calorimeter HV trips
    00:40 HV trip was reset and the run resumes
    1:00 stop run 2468, call for controlled access to change calorimeter angle
    3:40 Calorimeter in position, spectrometer angle and momentum have been set
    to -3.5 GeV, 14.58 degrees, target in LH2 no radiator position.
    3:45 Eugene begins his DVCS test runs
    6:00 Beam Off Maintainance period begins.